OpenStudy (anonymous):

square root of 5^8n = 125^(n + 3) ?

OpenStudy (saifoo.khan):

\[\Large 5^{8n} = 5^{3(n+3)}\]Can you solve.

OpenStudy (anonymous):

can only be a whole number

OpenStudy (anonymous):

that is the answer.

OpenStudy (anonymous):

no its not.

OpenStudy (saifoo.khan):

Try substituting the answer.

OpenStudy (anonymous):

Which of the following is the solution to the equation square root of 58n = 125(n + 3) ? Answer n = 3 n = 9 n = –9 n = –3

OpenStudy (anonymous):

since im obviously wrong

OpenStudy (anonymous):

yes it is wrong but how cuould take the square root of an equality?

OpenStudy (anonymous):

square root 5 answer power 8n =125^(n+3)

OpenStudy (anonymous):

i dont understand you

OpenStudy (anonymous):

women and math not ur best subject i guess

OpenStudy (saifoo.khan):

lol

OpenStudy (anonymous):

done.

OpenStudy (precal):